# maximum a posteriori

### v1

@@ Abstract
We introduce BOLT-GAN, a simple yet effective modification of the WGAN framework inspired by the Bayes optimal learning threshold (BOLT). We show that with a Lipschitz continuous discriminator, BOLT-GAN implicitly minimizes a different metric distance than the Earth-Mover (Wasserstein) distance, and also has better training stability. Empirical evaluations on four standard image-generation benchmarks, (CIFAR-10, CelebA-64, LSUN Bedroom-64, and LSUN Church‐64) show that BOLT-GAN consistently outperforms WGAN, achieving 10-60% lower Fréchet Inception Distance (FID). Our results suggest that BOLT is a broadly applicable principle for enhancing GAN training. The software is available in our anonymous repository.
@@ 1 Introduction
The Bayes optimal learning threshold (BOLT) (Naeini et al.,, 2025) is a recently proposed framework for discriminative learning. In its basic form, BOLT develops a computational approach to estimate the Bayes error rate, i.e., the lowest classification error achievable by a representation, by sampling an arbitrary learning model. This introduces the BOLT loss, whose minimization drives the model toward the Bayes optimal classifier.
This work adapts the BOLT framework to generative modeling by training the discriminator of a generative adversarial network (GAN) via the BOLT loss. We treat the GAN discriminator as a binary classifier and train it using the BOLT framework. Under this framework, the BOLT objective recovers the well-known Wasserstein GAN (WGAN) (Arjovsky et al.,, 2017) formulation, up to a minor adjustment: the discriminator outputs are constrained to a bounded range. This perspective provides both a Bayesian interpretation of adversarial learning and a pathway to improve training stability. Building on this foundation, we propose the BOLT-GAN, which inherits WGAN-style convergence guarantees while driving discriminator performance toward Bayes optimality.
@@ 1.1 Background and Related Work
The Bayes error rate is a central concept in statistical learning, characterizing the minimum error of a classification task (Bishop,, 2006). Classical approaches estimate this error rate via distance-based measures such as Bhattacharyya or Mahalanobis distances (Bhattacharyya,, 1943; Mahalanobis,, 1936). These methods often fail in high dimensions (Devroye et al.,, 1996). More recent methods employ nonparametric or model-based estimators, such as $k$-nearest neighbors (Moon et al.,, 2017) or $\epsilon$-ball techniques (Noshad et al.,, 2019). The BOLT framework (Naeini et al.,, 2025) provides a unifying alternative, deriving a bound on Bayes error from classifier outputs and introducing a loss that aligns learning with approximation of the Bayes optimal classifier.
@@
GAN (Goodfellow et al.,, 2014) consists of a generator, which synthesizes data samples from a latent variable, and a discriminator, which distinguishes between real and generated samples. It is trained via a min-max game between the generator and discriminator. While powerful, the original formulation based on Jensen–Shannon divergence is unstable (Goodfellow et al.,, 2014; Arjovsky et al.,, 2017). Numerous alternatives have been proposed: Wasserstein GAN (WGAN) (Arjovsky et al.,, 2017) leverages the Lipschitz continuity of the discriminator to improve gradient backpropagation. This is implemented by approaches such as WGAN-GP (Gulrajani et al.,, 2017) which enforce the Lipschitz constraint via gradient penalty. Other variants include $f$-GAN (Nowozin et al.,, 2016), which generalizes the objective to any $f$-divergence, LS-GAN (Mao et al.,, 2017), BEGAN (Berthelot et al.,, 2017), and techniques such as spectral normalization (Miyato et al.,, 2018) or progressive growing (Karras et al.,, 2018). These lines of work aim to stabilize training, but do not try to achieve statistical optimality.
Statistical and Bayesian Perspectives
The GAN has been studied from many statistical or Bayesian viewpoints. Bayesian GANs (Saatchi and Wilson,, 2017; Ye and Zhu,, 2018) introduce uncertainty quantification by estimating posterior distribution. Others, such as (Zhang et al.,, 2021; Arora et al.,, 2017), analyze GAN training through the lens of generalization and statistical distances. Our paper adds to this line of study by describing the GAN discriminator explicitly as a model-free Bayes-aligned classifier. In doing so, we provide a Bayesian perspective for adversarial learning and introduce a learning objective function, the bolt loss, for learning a representation that minimize the classification error.
@@ 1.2 Main Contributions
The key contributions of this work are as follows: ($i$) We introduce BOLT-GAN, an adversarial framework in which the discriminator is trained with the BOLT loss. The generator is implicitly trained to maximize the Bayes error of discrimination task. The inner maximization identifies the Bayes-aligned discriminator with the largest real-fake gap, and the outer minimization drives the generator to reduce this gap, effectively pushing the Bayes error toward random guessing. ($ii$) We characterize convergence of the BOLT-GAN. We show that, under balanced class priors, the BOLT objective is directly linked to total variation (TV) between the data and generator distributions. This finding shed lights on the numerical instability of the plain BOLT-GAN training algorithm. (iiiiii) To address this instability issue, we propose Lipschitz BOLT-GAN, which enforces a 1-Lipschitz constraint on the discriminator. We show that under this constraint, the learning objective is bounded above Wasserstein distance. This indicates that this framework guides the generator toward the data distribution using a weaker notion of distance. This leads to stabilization of the gradient backpropagation. ($iv$) We validate our results empirically. On CIFAR-10, CelebA-64, and LSUN, Lipschitz BOLT-GAN showing significant improvements.
@@
Throughout the paper, we adapt the following notation: expectations of of a random variable $X$ with respect to a distribution $P$ is denoted by $\mathbb{E}_{X\sim P}[\cdot]$. The distribution $P=P_{C}$ is defined as the conditional data distribution given the class C. The set $\left\{1,\ldots,n\right\}$ is abbreviated as $[n]$.
2 Background and Preliminaries
We restrict our analysis to a binary classification task with classes $C_{1}$ and $C_{2}$, although extensions to multi-class problems are possible. Consider a binary classification task with classes $C_{1}$ and $C_{2}$ occurring with prior probabilities $q_{1}$ and $q_{2}=1-q_{1}$, respectively. The data samples are drawn from the space $\mathcal{X}\subseteq\mathbb{R}^{d}$, for some integer $d$, with class-conditional distributions $P_{1}$ and $P_{2}$ for the classes $C_{1}$ and $C_{2}$, respectively. For this setting, the Bayes error rate (BER), denoted by $\varepsilon_{\mathrm{Bayes}}$, is defined as the minimum error probability achievable by a measurable classifier $f:\mathcal{X}\mapsto\{C_{1},C_{2}\}$. The BER is determined from the data distribution as (Bishop,, 2006)
$\varepsilon_{\mathrm{Bayes}}$ $=1-\int_{\mathcal{X}}\max_{\lambda}\left[q_{\lambda}P_{\lambda}(x)\right]\,\mathrm{d}x.$ (1)
and is achieved by the maximum a posteriori (MAP) classifier, which assigns a sample $x$ to the class $\lambda\in\{1,2\}$ that maximizes the joint distribution $q_{\lambda}P_{\lambda}(x)$.
2.1 BOLT Loss for Classification
An explicit computation of the BER(1) requires full knowledge of the data distribution, i.e.,  both the class prior $q_{\lambda}$ and the class conditional feature distribution $P_{\lambda}$ , As these are unavailable, the BER must be bounded and/or estimated ; see (Devroye et al.,, 1996; Fukunaga,, 2013; Noshad et al.,, 2019) for some sample studies. The recent study by Naeini et al., (2025) develops a new framework to estimate an upper bound on the Bayes error. The bound is universal in the sense that it can be computed by sampling from an arbitrary measurable and bounded function, e.g., a neural network. It is further tight, as it can achieve the Bayes error when optimized over a set of measurable functions. Motivated by these observations, Naeini et al., (2025) introduces the BOLT loss, which utilizes the proposed bound to approximate the optimal Bayes classifier via a parameterized learning model.
Theorem 1 (Binary Bayes error (Naeini et al.,, 2025)).
Let $h:\mathcal{X}\to[-1,0]$ be a measurable function. Then, the Bayes error of the binary classification satisfies
$\varepsilon_{\mathrm{Bayes}}\leq q_{2}+q_{2}\mathbb{E}_{X\sim P_{2}}\left[h(X)\right]-\hskip-1.42262ptq_{1}\mathbb{E}_{X\sim P_{1}}\left[h(X)\right].$ (2)
@@ 2.1 BOLT Loss for Classification
Naeini et al., (2025) show that the bound (2) is tight when optimized over $h$. This enables a computational means to estimate $\varepsilon_{\mathrm{Bayes}}$ by sampling from output of classifier. In particular if $h$ is a parameterized predictor model $h_{\theta}(\cdot)$, e.g., a neural network with weights $\theta$, the bound can be estimated and minimized. This transforms the upper bound in Theorem 1 into a training objective referred to as the Bayes optimal learning threshold (BOLT). While Naeini et al., (2025) establish the BOLT loss for a general multi-class problem, in this work we focus on the binary case, relevant to the discrimination component of the GANs.
Training via BOLT
Consider a training dataset $\mathcal{D}=\left\{(x_{i},\lambda_{i}):i\in[n]\right\}$, where $\lambda_{i}\in\{1,2\}$ denotes the class index from which $x_{i}$ is drawn. Assume that the samples in $\mathcal{D}$ are drawn i.i.d.  from the data distribution. BOLT trains a parameterized model $h_{\theta}:\mathcal{X}\rightarrow[-1,0]$ on the dataset $\mathcal{D}$ to estimate the Bayes error rate by optimizing BER empirical loss over $\theta$.
Remark 1.
Note that the output of the model $h_{\theta}$ is restricted to be within the interval [1,0][-1,0]. This is not a significant restriction, as it corresponds to a simple negation of the softmax output of the classifier.
For a given sample $(x,\lambda)$$\in\mathcal{D}$, the BOLT loss is defined as
$\ell_{\rm{BOLT}}(h_{\theta}(x),\lambda)=(-1)^{\lambda}h_{\theta}(x).$ (3)
To train the model $h_{\theta}$ on $\mathcal{D}$, where the empirical risk
$\hat{R}(\theta)=\frac{1}{n}\sum_{i\in[n]}\ell_{\text{BOLT}}(h_{\theta}(x_{i}),\lambda_{i}),$ (4)
This minimization can be done by mini-batch stochastic gradient descent or other standard training method.
The connection, between the BOLT-trained model and the Bayes error rate, can be understood using the results of Naeini et al., (2025): from Theorem 1, we can write
$\varepsilon_{\rm{Bayes}}\leq q_{2}+\mathbb{E}\left[\ell_{\rm{BOLT}}(h_{\theta}(X),\lambda)\right],$ (5)
where the expectation is taken with respect to the data distribution. Note that the prior $q_{2}$ is efficiently estimated from the data. By universal approximation (Cybenko,, 1989) and the tightness result of Naeini et al., (2025), one can conclude that the trained model in this case efficiently estimates the Bayes error.
@@ 2.2 Generative Adversarial Networks
A GAN consists of two learning models: a generator $g_{\phi}$ that generates data samples from the latent variable [^1] Typically, $Z$ is drawn from a multivariate Gaussian distribution. $Z\sim P_{Z}$ to generate samples of an estimator of $\mathcal{X}$, and a discriminator $h_{\theta}$ that distinguishes the generated sample $g_{\phi}(Z)$ from the true data sample $X\sim P_{\rm data}$. To keep the notation tractable, we denote the induced distribution, i.e., distribution of $g_{\phi}(Z)$, by $P_{\rm g}$, which is determined by $P_{Z}$ and the generator model $g_{\phi}$.
From a learning perspective, the discriminator solves a binary classification problem: it classifies the input as $C=$ real or $C=$ fake, corresponding to $\lambda=1$ or $\lambda=2$. The conventional approach is to train $h_{\theta}$ such that it estimates the posterior class probability. This leads to the vanilla GAN design proposed by Goodfellow et al., (2014). An alternative approach is to use the Kantorovich-Rubinstein duality (Villani et al.,, 2009) to implicitly minimize the Wasserstein distance between the data and induced distributions. This approach leads to the Wasserstein GAN design (Arjovsky et al.,, 2017). We briefly explain these two approaches in the sequel.
@@
The vanilla training approach, proposed in the initial study by Goodfellow et al., (2014), treats the discrimination as a classic binary classification problem. In this sense, it sets $h_{\theta}(x)=P(C=\texttt{real}|x)$ and uses the cross-entropy loss to train it. [^2] In practice, $h_{\theta}$ is a network with sigmoid-activated output. This approach implicitly implements the maximum likelihood training of the generator (Arjovsky et al.,, 2017). The risk function, i.e., the expected loss used for training, in this case is given by
$R_{\rm ML}(\phi,\theta)=\mathbb{E}\left[\mathrm{CE}(Y,\lambda)\right],$ (6)
where $\mathrm{CE}$ denotes the cross-entropy loss function, $Y$ is the output of the discriminator, i.e., $Y=h_{\theta}(X)$, for an input $X\in\mathcal{X}$ to the discriminator, and $\lambda\in\left\{1,2\right\}$ is the label of $X$ with $\lambda=1$ and $\lambda=2$ denoting real and fake inputs respectively. Where $\lambda$ is sampled from a prior $\pi=\Pr\left\{\lambda=1\right\}$ for some $\pi\in(0,1)$, [^3] Typically, $\pi=0.5$ is considered. and $X$ is sampled conditionally from $P_{1}=P_{\rm data}$ when $\lambda=1$ and $P_{2}=P_{\rm g}$ when $\lambda=2$.
Noting that $P_{2}=P_{\rm g}$ is the distribution of $g_{\phi}(Z)$ computed from the latent sample $Z\sim P_{Z}$, we can use the definition of cross-entropy to expand the risk (6) as
$R_{\rm ML}(\phi,\theta)=$ $-\pi\mathbb{E}_{X\sim P_{\rm{data}}}\left[\log h_{\theta}(X)\right]$ (7)
$-(1-\pi)\mathbb{E}_{Z\sim P_{Z}}\left[\log(1-h_{\theta}(g_{\phi}(Z)))\right].$
Defining $\mathcal{L}_{\rm ML}(\phi,\theta)=-R_{\rm ML}(\phi,\theta)$, the generator and discriminator are jointly trained by the following min-max game
$\min_{\phi}\max_{\theta}\mathcal{L}_{\rm ML}(\phi,\theta),$ (8)
which is equivalent to the worst-case risk minimization, i.e., fooling the best discriminator. It is shown that the solution to this min-max game, when solved for any measurable generator $g$ and discriminator $h$, guarantees the convergence of $P_{\rm g}$ to $P_{\rm data}$ as measured by the convergence of the Jensen–Shannon (JS) divergence to zero(Goodfellow et al.,, 2014, Theorem 1).
@@
Arjovsky et al., (2017) propose to train the generator-discriminator pair by minimizing the Wasserstein-$1$ distance, also known as Earth Mover (EM) distance, between $P_{\rm{data}}$ and $P_{\rm g}$. Using the Kantorovich-Rubinstein duality, it is shown that the EM distance minimization is equivalent to solving a min-max problem with the following objective (Arjovsky et al.,, 2017, Theorem 3)
$\mathcal{L}_{\rm EM}(\phi,\theta)=$ $\;\mathbb{E}_{X\sim P_{\rm{data}}}[h_{\theta}(X)]$
$-\mathbb{E}_{Z\sim P_{Z}}[h_{\theta}(g_{\phi}(Z))].$ (9)
@@
The training is carried out by solving the same min-max game as in (8) with $\mathcal{L}_{\rm EM}(\phi,\theta)$ as the objective. While the WGAN still guarantees the convergence of $P_{\rm g}$ to $P_{\rm data}$, the notion of convergence is weaker than vanilla GAN, in the sense that it only minimizes the EM distance between the two distributions instead of an information divergence. Arjovsky et al., (2017) show that the WGAN has better training stability, as compared with the vanilla GAN approach.
Remark 2.
Note that in the WGAN, the discriminator model $h_{\theta}$ is not restricted to compute probabilities. Intuitively, The discriminator in this case is a real-valued critic rather than a probabilistic classifier. The model $h_{\theta}$ is however constrained to be 1-Lipschitz. In practice, this constraint can be enforced by approaches such as weight clipping, gradient penalty, or spectral normalization. See the supplementary file for additional discussion.
3 BOLT-GAN Framework
Here we develop the proposed BOLT-GAN framework for training a GAN model. This scheme invokes the BOLT loss to train the generator to fool the Bayes optimal discriminator. The convergence in BOLT-GAN is in the sense of convergence in total variation, i.e., stronger than vanilla GAN, for a Lipschitz discriminator it converges in a weaker sense than WGAN. This implies a trade-off between the strength of convergence and stability.
3.1 MAP Estimator via BOLT
To understand the BOLT-GAN framework, we first present a result which relates a BOLT-trained model to the Bayesian optimal classifier: consider $h_{\theta}:\mathcal{X}\rightarrow[-1,0]$ that is trained for a binary classification task using the BOLT loss. Theorem 2 shows how an estimator of the Bayesian optimal classifier, i.e., the MAP estimator, is computed from $h_{\theta}$.
Theorem 2 (BOLT Estimator of MAP Classifier).
Let $h^{\star}$ be the minimizer of the upper bound in Theorem 1. Assume that the model $h_{\theta}$ approximates $h^{\star}$ in the sense that $\|h^{\star}(x)-h_{\theta}(x)\|_{\infty}\leq\epsilon$. Then, the plug-in classifier
$\widehat{C}_{\theta}(x)=\begin{cases}C_{1}&h_{\theta}(x)\geq-0.5,\\[4.0pt] C_{2}&h_{\theta}(x)<-0.5.\end{cases}$ (10)
is a plug-in estimator of the MAP classifier, i.e.
$\widehat{C}(x)=\arg\max_{\lambda\in\{1,2\}}q_{\lambda}P_{\lambda}(x)$ (11)

### v2

@@ Abstract
We introduce BOLT-GAN, a novel framework for stable GAN training using the Bayes optimal learning threshold (BOLT). The discriminator is trained via the BOLT loss under a standard 1-Lipschitz constraint. This guides the generator to maximize the Bayes error of the discrimination task. We show that the training objective in this case represents a class of metrics on probability measures controlled by a 1-Lipschitz discriminator minimizing an integral probability metric that is upper-bounded by Wasserstein-1 distance. Across four standard image-generation benchmarks, BOLT-GAN improves FID and precision/recall over benchmark GAN frameworks under identical architectures and training budgets. Our experimental findings further confirm the advantage of linking the GAN training objective to a min-max Bayes error criterion. Implementation software is available in our anonymous repository.
@@ 1 Introduction
The Bayes optimal learning threshold (BOLT) (Naeini et al., 2025) is a recently proposed framework for discriminative learning. In its basic form, BOLT provides a computational approach to upper-bound (and tightly approximate) the Bayes error rate (BER) by sampling an arbitrary bounded learning model. Minimizing the resulting BOLT loss, aligns the model with a Bayes optimal, i.e., maximum-a-posteriori (MAP), classifier.
Generative adversarial networks (GANs) learn data distribution through a two-player game between a generator and a discriminator. Despite their success, vanilla GANs trained with the cross-entropy loss and Jensen-Shannon divergence are notoriously unstable, often suffering from vanishing gradients and mode collapse. Since the discriminator’s role is inherently a binary classification task (real versus fake), it is natural to ask whether a Bayes-error-motivated discriminator objective can improve generative training dynamics. In this work, we embed the BOLT loss inside the adversarial game: we train the discriminator with the BOLT objective and train the generator to maximize the induced Bayes error rate of the discrimination task. We refer to this adversarial formulation as a maximum Bayes error (Max-BER) game.
Our analysis further shows that the Max-BER game induces a discrepancy between the data distribution and the generator distribution that is controlled by the discriminator. With unregularized bounded-discriminator, this game reduces under balanced priors to minimizing the total variation (TV), a strong metric. This can yield poor gradient propagation and severe numerical instability. By restricting the discriminator class to bounded 1-Lipschitz functions, we obtain a weaker but smoother metric: an integral probability metric (IPM) that is upper-bounded by the Wasserstein-1 distance. We refer to this constrained min-max formulation as BOLT-GAN. Unlike existing GAN formulations, BOLT-GAN provides both a Bayesian interpretation of adversarial training and a pathway to improved stability.
@@ 1.1 Background and Related Work
The Bayes error rate (BER) characterizes the minimum achievable classification error for a given data-generating process (Bishop, 2006). Classical BER bounds and estimators are often expressed via distance-like quantities between class-conditional distributions (e.g., Bhattacharyya- or Mahalanobis-type measures), but they can be inaccurate in high dimensions and under model mismatch (Devroye et al., 1996). Modern nonparametric approaches estimate BER using nearest-neighbor and local neighborhood constructions (Moon et al., 2017; Noshad et al., 2019). The Bayes optimal learning threshold (BOLT) framework (Naeini et al., 2025) provides a universal upper bound on BER based on bounded discriminator outputs and induces a surrogate loss (the BOLT loss) whose minimization aligns a model with the Bayes-optimal, i.e. maximum-a-posteriori (MAP), classifier.
@@
A GAN trains a generator and a discriminator via a min-max game (Goodfellow et al., 2014). The original GAN training objective corresponds to optimizing the Jensen-Shannon (JS) divergence between the generative and data distributions. This basic approach is known to suffer from gradient saturation and instability, especially when the model and data supports have limited overlap (Goodfellow et al., 2014; Arjovsky et al., 2017). Alternatives replace JS with other metrics and discrepancies, e.g. $f$-divergence (Nowozin et al., 2016), least-squares losses (Mao et al., 2017), and auto-encoding reconstruction-loss (Berthelot et al., 2017). Among these designs, Wasserstein GAN (WGAN) shows substantial training stability (Arjovsky et al., 2017). WGAN replaces JS divergence with the Wasserstein-1 distance $W_{1}$, yielding a smoother training signal and improved stability.
Stable GAN Training
A central component in WGAN training is enforcing the 1-Lipschitz constraint on the discriminator function. This is commonly addressed via gradient penalty (Gulrajani et al., 2017) or spectral normalization (Miyato et al., 2018). Other stabilization approaches include training heuristics such as progressive growing (Karras et al., 2018). Bayesian GAN formulations place priors over network parameters and sample from posterior distributions to quantify uncertainty (Saatchi and Wilson, 2017; Ye and Zhu, 2018). Several studies analyze generalization and equilibrium behavior via statistical distances for learning dynamics (Arora et al., 2017; Zhang et al., 2021). Our work complements these lines of work by directly training the discriminator for Bayes optimal discrimination via the BOLT loss.
@@ 1.2 Main Contributions
The key contributions of this work are as follows. (i) Bayes-optimal GAN training via Max-BER: we deploy the BOLT framework to formulate Bayes-optimal GAN training, where the generator is trained to maximize the Bayes error rate of the discrimination task. We first show the Bayes optimality of BOLT (Theorem 2) and characterize the bias and variance of the BOLT plug-in estimation of the BER (Theorem 3). We then conclude that under the Max-BER game, the generator and data distributions match in total variation (Theorem 4). (ii) BOLT-GAN framework: We further show that, when the discriminator is restricted to the class of Lipschitz models, the BOLT-trained GAN minimizes the distance between the generator and data distributions with respect to an alternative IPM that is bounded by the Wasserstein-1 distance (Theorem 5). The resulting BOLT-GAN trains the generator toward the true data distribution using a weaker notion of distance, leading to more efficient gradient backpropagation and hence better stability. (iii) Empirical validation: We validated BOLT-GAN through experiments with the CIFAR-10, CelebA-64, LSUN Bedroom-64, and LSUN Church-64 datasets. We report FID, precision, and recall, and compare against WGAN and HINGE-GAN baselines (Miyato et al., 2018).
@@
On a measurable space $(\mathcal{X},\mathcal{F})$, probability measures are shown uppercase, e.g., $P$, and their corresponding densities are denoted lowercase, e.g., $p$. Expectation with respect to probability measure $P$ is shown by $\mathbb{E}_{X\sim P}[\cdot]$. The set $\{1,\dots,n\}$ is abbreviated as $[n]$.
2 Preliminaries
We consider binary classification with classes $C\in\{1,2\}$ and data space $\mathcal{X}\subseteq\mathbb{R}^{d}$. Let $q_{i}:=\Pr(C=i)$ denote the class priors ($q_{2}=1-q_{1}$), and let $P_{i}$ denote the class-conditional distribution of $X$ given $C=i$. When needed, assume $P_{i}$ admits a density $p_{i}$ with respect to a common dominating measure $\mu$ on $\mathcal{X}$.
The Bayes error rate (BER), denoted by $\varepsilon_{\rm Bayes}$, is the minimum error probability achievable by any measurable classifier $f:\mathcal{X}\to\{C_{1},C_{2}\}$, which is achieved by the maximum a posteriori (MAP) rule (Bishop, 2006). In terms of densities, the BER is
$\varepsilon_{\rm Bayes}=1-\int_{\mathcal{X}}\max_{i\in\{1,2\}}\left\{q_{i}\,p_{i}(x)\right\}\,d\mu(x),$ (1)
and the MAP classifier assigns sample $x$ to
$\widehat{C}(x)\in\operatorname*{arg\,max}_{i\in\{1,2\}}q_{i}\,p_{i}(x).$ (2)
2.1 BOLT Framework
With unknown data distribution, the BER can be approximated (Devroye et al., 1996; Fukunaga, 2013; Noshad et al., 2019). Naeini et al. (2025) have developed an empirical estimate of universal upper bound on BER that can be computed by sampling an arbitrary measurable and bounded function (e.g., a neural network). The resulting surrogate objective given by this upper bound is called the BOLT risk.
The following theorem presents the universal bound for binary classification.
Theorem 1 (Binary Bayes error (Naeini et al., 2025)).
Let $h:\mathcal{X}\to[0,1]$ be a measurable function. Then, the BER is bounded as  [^1] The result in Naeini et al. (2025) is given for $\tilde{h}:\mathcal{X}\to[-1,0]$. We represent the result after the reparameterization $h=1+\tilde{h}$.
$\varepsilon_{\rm Bayes}\leq q_{1}+q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]-q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)].$ (3)
@@ 2.1 BOLT Framework
Naeini et al. (2025) show that the bound (3) is tight when optimized over $h$. This enables a computational means to estimate $\varepsilon_{\rm Bayes}$ by sampling the output of a bounded computational model $h$ with parameters $\theta$. This transforms the upper bound in Theorem 1 into a training objective referred to as the Bayes optimal learning threshold (BOLT). While Naeini et al. (2025) establish the BOLT loss for a general multi-class problem, we focus on the binary case relevant to the discrimination component of GANs.
BOLT loss
Considering a labeled dataset $\mathcal{D}=\{(x_{i},C_{i}):i\in[n]\}$, where $C_{i}\in\{1,2\}$. BOLT trains the model $h_{\theta}:\mathcal{X}\to[0,1]$ by minimizing an empirical surrogate of the bound in Theorem 1. This can be seen as an empirical risk minimization problem, where each $(x,C)\in\mathcal{D}$ is a (feature,label) sample and the BOLT loss is given by
$\ell_{\rm BOLT}(h_{\theta}(x),C)=(-1)^{C}\,h_{\theta}(x).$ (4)
The bound in Theorem 1 can be directly written in terms of BOLT risk as
$\varepsilon_{\rm Bayes}\leq q_{1}+\mathbb{E}\bigl[\ell_{\rm BOLT}(h_{\theta}(X),C)\bigr],$ (5)
where the expectation is over $(X,C)$ generated as $\Pr(C=i)=q_{i}$ and $X\mid\{C=i\}\sim P_{i}$, for $i\in\{1,2\}$.
@@ 2.2 Generative Adversarial Networks
A GAN consists of two functions: a generator $g:\mathcal{Z}\to\mathcal{X}$ that maps a latent variable $Z\sim P_{Z}$ to a sample in data space $\mathcal{X}$, and a discriminator $h:\mathcal{X}\to[0,1]$ that distinguishes generated samples $g(Z)$ from true data samples $X\sim P_{\rm data}$. We denote the induced distribution of $g(Z)$ by $P_{\rm g}$.
@@
The vanilla training approach, proposed in the initial study by Goodfellow et al. (2014), treats the discrimination as a classic binary classification problem. In this sense, it sets $h(x)=P(C=\texttt{real}|x)$ and uses the cross-entropy loss to train it. This approach implicitly implements the maximum likelihood training of the generator (Arjovsky et al., 2017). The risk function, i.e., the expected loss used for training, in this case is given by
$R_{\rm ML}(g,h)=\mathbb{E}\left[\mathrm{CE}(Y,C)\right],$ (6)
where $\mathrm{CE}$ denotes the cross-entropy loss function, $Y$ is the output of the discriminator, i.e., $Y=h(X)$, for an input $X\in\mathcal{X}$ to the discriminator, and $C\in\left\{1,2\right\}$ is the label of $X$ with $C=1$ and $C=2$ denoting real and fake inputs respectively. Where $C$ is sampled from a prior $\pi=\Pr\left\{C=1\right\}$ for some $\pi\in(0,1)$, [^2] Typically, $\pi=0.5$ is considered. and $X$ is sampled conditionally from $P_{1}=P_{\rm data}$ and $P_{2}=P_{\rm g}$.
Noting that $P_{2}=P_{\rm g}$ is the distribution of $g(Z)$ computed from the latent sample $Z\sim P_{Z}$, we can use the definition of cross-entropy to expand the risk in (6) as
$R_{\rm ML}(g,h)=$ $-\pi\mathbb{E}_{X\sim P_{\rm{data}}}\left[\log h(X)\right]$ (7)
$-(1-\pi)\mathbb{E}_{Z\sim P_{Z}}\left[\log(1-h(g(Z)))\right].$
Defining $\mathcal{L}_{\rm ML}(g,h)=-R_{\rm ML}(g,h)$, the generator learns to sample $P_{\rm data}$ by the following min-max game
$\min_{g}\max_{h}\;\mathcal{L}_{\rm ML}(g,h),$ (8)
which is equivalent to  fooling the best discriminator. It is shown that the solution to this min-max game, when solved for any measurable generator $g$ and discriminator $h$, guarantees the convergence of $P_{\rm g}$ to $P_{\rm data}$ as measured by the convergence of the Jensen–Shannon (JS) divergence to zero (Goodfellow et al., 2014, Theorem 1).
@@
Arjovsky et al. (2017) propose to train the generator–discriminator pair by minimizing the Wasserstein-$1$ distance (Earth Mover distance) between $P_{\rm data}$ and $P_{\rm g}$. Using the Kantorovich–Rubinstein duality, this distance can be written as a supremum over $1$-Lipschitz functions; accordingly, the WGAN “discriminator” is a real-valued critic (rather than a probabilistic classifier) and is constrained to be $1$-Lipschitz in theory. In practice, this constraint is typically enforced via weight clipping, gradient penalty, or spectral normalization (see the supplementary material for additional discussion). This yields the min–max objective (Arjovsky et al., 2017, Theorem 3)
$\mathcal{L}_{\rm EM}(g,h)=\mathbb{E}_{X\sim P_{\rm data}}[h(X)]-\mathbb{E}_{Z\sim P_{Z}}[h(g(Z))].$ (9)
@@
Learning proceeds by solving the same min-max game as in (8) with $\mathcal{L}_{\rm EM}(g,h)$ as the objective. While WGAN still guarantees convergence of $P_{\rm g}$ to $P_{\rm data}$, the notion of convergence is in Wasserstein distance (a weaker topology than that induced by common information divergences), which often leads to improved training stability compared with vanilla GAN (Arjovsky et al., 2017).
3 GAN with Maximum BER
Discriminator training in GANs is inherently a binary classification problem (real versus fake), which motivates using BOLT. However, deploying BOLT inside the adversarial game requires understanding how the Bayes-optimal discriminator is estimated from a BOLT-trained model. In the sequel, we characterize this connection by showing that a population BOLT minimizer recovers the Bayes optimal discriminator by thresholding. We further derive a finite-sample guarantee for a plug-in BER estimator computed by empirical BOLT risk. These results enable us to formulate the GAN training via a min-max problem, in which the generator is trained adversarially to increase the discriminator-induced BER. We characterize the metric under which $P_{\rm g}$ converges to $P_{\rm data}$.
3.1 Bayes Optimal Discriminator via BOLT
Naeini et al. (2025) conjectured that a model trained via empirical BOLT risk can estimate the Bayes optimal classifier, and proposed a heuristic plug-in rule based on thresholding the learned score $h$. In the sequel, we formalize this connection. Theorem 2 shows that any population minimizer of the BOLT risk is Bayes-aligned and that thresholding at $1/2$ recovers the MAP decision rule.
Theorem 2 (MAP vs. BOLT optimum).
Consider the binary classification setup in Section 2 and let $P_{X}$ denote the marginal distribution of $X$. Let
$h^{\star}\in\operatorname*{arg\,min}_{h:\mathcal{X}\to[0,1]}\;\mathbb{E}\bigl[\ell_{\rm BOLT}(h(X),C)\bigr].$ (10)
and
$\eta(x):=\Pr\left(C=1\mid X=x\right)$ (11)
be the likelihood of $x$. Then, for $P_{X}$-a.e. $x$, we have
$h^{\star}(x)=\begin{cases}1,&\eta(x)>\tfrac{1}{2},\\ 0,&\eta(x)<\tfrac{1}{2},\\ \text{z},&\eta(x)=\tfrac{1}{2}.\end{cases}$ (12)
for some $z\in[0,1]$. Consequently, the plug-in classifier
$\widehat{C}(x)=\begin{cases}1,&h^{\star}(x)\geq\tfrac{1}{2},\\ 2,&h^{\star}(x)<\tfrac{1}{2},\end{cases}$ (13)
coincides $P_{X}$-a.e. with the MAP classifier
$C_{\rm MAP}(x)\in\operatorname*{arg\,max}_{k\in\{1,2\}}\Pr\{C=k\mid X=x\}.$ (14)